OpenStudy (anonymous):

Find the least common multiple (LCM) (3+2P), (9-4p^2), (3-2p)

OpenStudy (anonymous):

(9-4P^2) or (3+2P)(3-2P)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

oh ok here

OpenStudy (anonymous):

you need the to factor \[9-4p^2=(3+2p)(3-2p)\]

OpenStudy (anonymous):

so your least common multiple is \[(3+2p)(3-2p)=9-4p^2\]